gianarapolla1:

a married couple had a combined annual income of 76,000. the wife made $4,000 more than her husband. what was each of their incomes

cyanide:

Let the husband's income be 'x'. x + x + 4,000 = 76,000 2x = 76,000 - 4,000 = 72,000 x = 36,000 (husband) x + 4,000 = 40,000 (wife)